# An Extension of the Adiabatic Theorem

### v1

@@ III Transverse Field Ising Model
This condition (17) is fulfilled if and only if the argument of $\tan^{2}(x)$ lies in the interval $[-\frac{\pi}{4},\frac{\pi}{4}]$. We conclude
$\bigg{|}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{|}\overset{!}{<}\frac{\pi}{4}.$ (18)
The angles $\theta_{k}(h)$ are known to lie in $[0,\frac{\pi}{2}]$ with their explicit expressions given by the parameterisation of the Bogoliubov transformation [11]
$\tan(2\theta_{k})=\frac{\sin k}{h-\cos k}\ \Leftrightarrow\ \theta_{k}(h)=\frac{1}{2}\arctan\bigg{(}\frac{\sin k}{h-\cos k}\bigg{)}.$ (19)
As $k=2\pi\frac{n}{L}$, with $n\in\mathbb{N}$ in the interval $[1,\frac{L}{2}]$ and $L$ being the spin chain length, the values for $k$ cover half a period from 0 to $\pi$. Using the inverse tangent function
$\tan^{-1}(2\theta_{k})=\begin{cases}\arctan(2\theta_{k}),\quad 2\theta_{k}>0\\ \arctan(2\theta_{k})+\pi,\quad 2\theta_{k}<0,\end{cases}$ (20)
implies for the definition of $\theta_{k}(h)$
$\theta_{k}(h)=\begin{cases}\frac{1}{2}\arctan\bigg{(}\frac{\sin k}{h-\cos k}\bigg{)},\quad\frac{\sin k}{h-\cos k}>0\\ \frac{1}{2}\bigg{(}\arctan(\frac{\sin k}{h-\cos k})+\pi\bigg{)},\quad\frac{\sin k}{h-\cos k}<0.\end{cases}$ (21)
  1. (i):
    Both Hamiltonians lie in the paramagnetic phase. Without loss of generality, it is assumed that $1<h_{i}<h_{f}$. In the paramagnetic phase, $\theta_{k}=\frac{1}{2}\arctan(\frac{\sin k}{h-\cos k})\forall h_{i},h_{f},k$ as $\frac{\sin k}{h-\cos k}>0$ for all values of $h_{i},h_{f}\in(1,\infty)$ and $k\in[0,\pi]$.
    The $\arctan(\frac{\sin k}{h_{i}-\cos k})$ is a monotonically decreasing function in $h$ for all $k\in[0,\pi]$. Therefore, the maximal difference of $\theta_{k}(h_{i})$ and $\theta_{k}(h_{f})$ is reached for maximally different $h_{i}$ and $h_{f}$. We choose $h_{i}$ to be $1+\epsilon$ with $\epsilon\rightarrow 0$ and $h_{f}$ to be $x\rightarrow\infty$ and obtain
    $\max_{h_{i},h_{f},k}\bigg{(}\bigg{|}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{|}\bigg{)}=\frac{\pi}{4}.$ (22)
    As the function approaches the limit from below, the conjecture (2) is fulfilled for all possible parameter combinations of $h_{i}$, $h_{f}$ and $k$ in the paramagnetic phase.
  2. (ii):
    Both Hamiltonians lie in the ferromagnetic phase. Without loss of generality, it is assumed that $10\leq h_{i}<h_{f}<$ 1. For $h\in[0,1)$, we rewrite Eq. (21):
    $\theta_{k}(h)=\begin{cases}\frac{1}{2}\arctan\bigg{(}\frac{\sin k}{h-\cos k}\bigg{)},\quad k>\arccos(h)\\ \frac{1}{2}\bigg{(}\arctan\bigg{(}\frac{\sin k}{h-\cos k}\bigg{)}+\pi\bigg{)},\quad k<\arccos(h).\end{cases}$ (23)
    We consider $\arccos(h)<k=\arccos(h)+\nu$ with $\nu>0$.
    $\begin{split}&\max_{h_{i},h_{f},k}\bigg{(}\bigg{|}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{|}\bigg{)}\\ =&\max_{h_{i},h_{f},\nu}\bigg{(}\frac{1}{2}\bigg{|}\arctan\bigg{(}\frac{\sin(\arccos(h_{i})+\nu)}{h_{i}-\cos(\arccos(h_{i})+\nu)}\bigg{)}\\ &\quad\quad\quad\ -\arctan\bigg{(}\frac{\sin(\arccos(h_{f})+\nu)}{h_{f}-\cos(\arccos(h_{f})+\nu)}\bigg{)}\bigg{|}\bigg{)}\end{split}$ (24)
    As the $\arctan(\cdot)$ is a monotone function in $h$ for all $\nu\in(0,\pi]$, the maximal difference of $\theta_{k}(h_{i})$ and $\theta_{k}(h_{f})$ is reached for maximally different $h_{i}$ and $h_{f}$. Thus, we choose $h_{i}$ to be 00 and $h_{f}$ to be $1-\epsilon$ with $\epsilon\rightarrow 0$, such that
    $=\max_{\nu}\bigg{(}$ $\frac{1}{2}\bigg{|}\arctan\bigg{(}\frac{\sin(\frac{\pi}{2}+\nu)}{\cos(\frac{\pi}{2}+\nu)}\bigg{)}$ (25)
    $+\arctan\bigg{(}\frac{\sin(\nu)}{1-\cos(\nu)}\bigg{)}\bigg{|}\bigg{)}$ (26)
    $\max_{h_{i},h_{f},k}\bigg{(}\bigg{|}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{|}\bigg{)}=\frac{\pi}{8}<\frac{\pi}{4}.$ (27)
  3. (iii):
    Again, both Hamiltonians lie in the ferromagnetic phase. As before, without loss of generality, it is assumed that $10<h_{i}<h_{f}<$ 1. However, now we consider the case of $\arccos(h)>k=\arccos(h)-\nu$ with $\nu>0$. From $k\in[0,\pi]$ it follows that $\nu\overset{!}{\in}[-\frac{\pi}{2},0)$.
    $\begin{split}&\max_{h_{i},h_{f},k}\bigg{(}\bigg{|}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{|}\bigg{)}\\ =&\max_{h_{i},h_{f},\nu}\bigg{(}\frac{1}{2}\bigg{|}\arctan\bigg{(}\frac{\sin(\arccos(h_{i})-\nu)}{h_{i}-\cos(\arccos(h_{i})-\nu)}\bigg{)}+\pi\\ &\quad\quad\quad\quad-\arctan\bigg{(}\frac{\sin(\arccos(h_{f})-\nu)}{h_{f}-\cos(\arccos(h_{f})-\nu)}\bigg{)}+\pi\bigg{|}\bigg{)}\end{split}$ (28)
    As $\arctan(\cdot)$ is a monotone function in $h$ for all $\nu\in[-\frac{\pi}{2},0)$, the maximal difference of $\theta_{k}(h_{i})$ and $\theta_{k}(h_{f})$ is again reached for maximally different $h_{i}$ and $h_{f}$. Thus we again choose $h_{i}$ to be 00 and $h_{f}$ to be $1-\epsilon$ with $\epsilon\rightarrow 0$, such that
    $\max_{h_{i},h_{f},k}\bigg{(}\bigg{|}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{|}\bigg{)}=\frac{\pi}{8}<\frac{\pi}{4},$ (29)
    and the conjecture (2) is confirmed for this second choice of $\theta_{k}(h)$.
However, these have even smaller overlap with the pre-quench ground state than the excitation (11), as each excitation pair contributes another factor further diminishing the overlap with the pre-quench ground state.
IV Axial Next Nearest Neighbour Ising Model
Figure 1: Phase diagram of the ANNNI model. The red line within the paramagnetic phase indicates the Peschel-Emery line. Adapted from [14].
The second model investigated in this work is a modification of the Transverse Field Ising Model, incorporating additional next nearest neighbour interactions. It is known as the Anisotropic or Axial Next Nearest Neighbour Ising Model and has been studied extensively in the last few decades [15, 16, 14, 7, 8]. Alongside the nearest neighbour coupling $J$ and the coupling $h$ to an external magnetic field, the interaction between next nearest neighbours can be tuned by a parameter $\Delta$ in the Hamiltonian [14]
@@ IV Axial Next Nearest Neighbour Ising Model
The competition between interactions gives rise to a more complex phase diagram, accommodating four different phases, as illustrated in Fig. 1. These phases are the paramagnetic (PM) phase, the ferromagnetic (FM) phase, a critical incommensurate floating phase (FP), and an anti-phase (AP).
In contrast to the TFIM the ANNNI model is not integrable [14], except along two lines in the phase diagram: the trivial line with $\Delta=0$ (that corresponds to the TFIM) and the so-called Peschel-Emery (PE) line [18]. The PE line is a disorder line along which the Hamiltonian factorises into local Hamiltonians $H_{j}^{\operatorname{loc}}$, such that it takes the form [19]
$H_{\text{ANNNI}}^{\text{PE}}=\sum_{j}H_{j}^{\operatorname{loc}}.$ (32)
The ground state of this factorised Hamiltonian is exactly degenerate and can be written as product states [20]. The model described by Eq. (32) is frustration-free, as the ground state of the full Hamiltonian $H_{\text{ANNNI}}^{\text{PE}}$ minimises each local Hamiltonian $H_{j}^{\operatorname{loc}}$ independently [19]. In the ANNNI model, the PE line is described by [18]
@@
The Hamiltonian of the ANNNI model can be mapped onto an interacting quantum Ising model by dual mapping and a subsequent rotation, resulting in the interacting quantum Ising Hamiltonian [20]. After further applying a Jordan-Wigner transformation, this Hamiltonian is equivalent to the Kitaev-Hubbard chain and the $XZ$ Heisenberg chain [20, 19, 21]:
$\begin{split}H_{\text{ANNNI}}=&J\sum_{j=1}^{L}\bigg{(}h\left(c_{j}^{\dagger}-c_{j}\right)\left(c_{j+1}^{\dagger}+c_{j+1}\right)\\ &+\Delta\left(1-2c_{j}^{\dagger}c_{j}\right)\left(1-2c_{j+1}^{\dagger}c_{j+1}\right)-\left(1-2c_{j}^{\dagger}c_{j}\right)\bigg{)}.\end{split}$ (34)
Due to the quartic interaction arising from the next-nearest neighbour coupling in the spin-$\frac{1}{2}$ system, the ANNNI model does not have a simple analytical solution like the TFIM.
The ground state along the PE line is a direct product state [19]:
$|\text{GS}_{i}^{(\pm)}\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}\frac{1}{\sqrt{(1\pm\alpha)^{L}}}\bigg{(}\left|\uparrow\right\rangle_{j}\pm\alpha\left|\downarrow\right\rangle_{j}\bigg{)},$ (35)
where $\alpha=\sqrt{\cot{\left(\frac{\theta^{*}}{2}\right)}}$ with $\theta^{*}=\arctan{\left(h^{*}\right)}=\arctan{\left(\Delta-\frac{1}{4\Delta}\right)}$ [19]. We see that $\alpha\geq 1$ along the entire frustration-free line.
As a post-quench Hamiltonian we consider the model at $\mu=0$,
$H_{0}=-t\sum_{j=1}^{L-1}\sigma_{j}^{x}\sigma_{j+1}^{x}.$ (36)
The eigenstates of this Hamiltonian are simply products of the eigenstates of $\sigma_{j}^{x}$, namely
$|\phi\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}|\pm\rangle_{j},$ (37)
and the ground states are
$|+\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}|+\rangle_{j}\quad\text{and}\quad|-\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}|-\rangle_{j}.$ (38)
Expressing $|\text{GS}_{i}^{(\pm)}\rangle$ in terms of |±$|\pm\rangle_{j}$ yields
$|\text{GS}_{i}^{(\pm)}\rangle=\frac{1}{\sqrt{2^{L}}}\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}\bigg{(}(1\pm\alpha)|+\rangle_{j}+(1\mp\alpha)|-\rangle_{j}\bigg{)},$ (39)
where we used
$\left|\uparrow\right\rangle_{j}$ $=\frac{1}{\sqrt{2}}\bigg{(}|+\rangle_{j}+|-\rangle_{j}\bigg{)}\quad\text{and}$ (40)
$\left|\downarrow\right\rangle_{j}$ $=\frac{1}{\sqrt{2}}\bigg{(}|+\rangle_{j}-|-\rangle_{j}\bigg{)}.$ (41)
The post-quench ground states are given by $|\phi_{0}^{+}\rangle=|\phi^{L}\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}|+\rangle_{j}$ and $|\phi_{0}^{-}\rangle=|\phi^{0}\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}|-\rangle_{j}$, whereas the post-quench excited states are given by $|\phi^{N}\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{N}}|+\rangle_{j}\mathbin{\mathop{\otimes}\displaylimits_{l=N+1}^{L}}|-\rangle_{l}$ and its permutations, where $N<L$. The resulting overlaps with the pre-quench ground state are
$\bigg{|}\langle\text{GS}_{i}^{(\pm)}|\phi_{0}^{\pm}\rangle\bigg{|}^{2}$ $=\frac{1}{2^{L}}\bigg{(}1\pm\alpha\bigg{)}^{2L}$ (42)
for the ground state overlap, and
$\bigg{|}\langle\text{GS}_{i}^{(\pm)}|\phi^{N}\rangle\bigg{|}^{2}$ $=\frac{1}{2^{L}}(1\pm\alpha)^{2N}(1\mp\alpha)^{2(L-N)}$ (43)
for the overlap of the pre-quench ground state with a post-quench excited state.
$\bigg{|}\langle\text{GS}_{i}|\text{GS}_{f}\rangle\bigg{|}^{2}\overset{!}{>}\bigg{|}\langle\text{GS}_{i}|\tilde{\psi}\rangle\bigg{|}^{2}\Leftrightarrow\bigg{|}\frac{\langle\text{GS}_{i}|\tilde{\psi}\rangle}{\langle\text{GS}_{i}|\text{GS}_{f}\rangle}\bigg{|}^{2}\overset{!}{<}1.$ (44)
We consider both ground states separately:
$\begin{split}&\bigg{|}\frac{\langle\text{GS}_{i}^{+}|\phi^{N}\rangle}{\langle\text{GS}_{i}^{+}|+\rangle}\bigg{|}^{2}=\frac{(1+\alpha)^{2N}(1-\alpha)^{2(L-N)}}{(1+\alpha)^{2L}}\\ =&\frac{(1-\alpha)^{2(L-N)}}{(1+\alpha)^{2(L-N)}}=\bigg{(}\frac{1-\alpha}{1+\alpha}\bigg{)}^{2(L-N)}<1\forall\alpha_{i}>0\end{split}$ (45)
$\begin{split}&\bigg{|}\frac{\langle\text{GS}_{i}^{-}|\phi^{N}\rangle}{\langle\text{GS}_{i}^{-}|-\rangle}\bigg{|}^{2}=\frac{(1-\alpha)^{2N}(1+\alpha)^{2(L-N)}}{(1+\alpha)^{2L}}\\ =&\frac{(1-\alpha)^{2N}}{(1+\alpha)^{2(L-(L-N))}}=\bigg{(}\frac{1-\alpha}{1+\alpha}\bigg{)}^{2N}<1\forall\alpha_{i}>0\end{split}$ (46)
[…]
Equations (45) and (46) confirm that condition (44) is satisfied. Thus, the conjecture could be verified analytically for the special case of a quench originating along the frustration-free PE line and ending at the specific choice of the Hamiltonian given in Eq. (36).
V Numerical Analysis of the ANNNI model
To investigate the conjecture’s validity beyond this special case, we conducted a numerical analysis of the ANNNI model. We present results obtained through Exact Diagonalisation (ED). Given the constraints of the available computational resources, the largest accessible system size for these computations is $L=16$ spins. The presentation of the results is structured according to the phase diagram shown in Fig. 1. Each point in the phase diagram corresponds to a parameter combination that uniquely defines the post-quench Hamiltonian. For clarity and ease of interpretation, dashed lines are used to reconstruct the phase boundaries. The quench starting points are indicated in the title of each plot and marked by a star in the phase diagram. The results are displayed using a colour code based on the normalised deviation of the ground state overlap from the maximal overlap, defined as:
$\text{deviation}=\frac{|\text{GS}\rangle\ \text{overlap}-\text{max. overlap}}{\text{max. overlap}}.$ (47)
In this visualisation, a very bright teal colour indicates that the ground state overlap equals the maximal overlap, while darker colours signify a greater deviation between these two quantities.
Within the phase where the quench starts, all points should ideally be displayed in the brightest colour for the conjecture to hold true. Any darker colour within that phase indicates a violation of the conjecture, as it signifies that the maximal overlap is not provided by the ground state overlap in such cases.
@@ V Numerical Analysis of the ANNNI model
Figure 2: Quenches within the ferromagnetic (a), floating (b) and anti-phase (c).
For all tested quench configurations within the ferromagnetic, floating and anti-phase, the numerical data consistently support the conjecture. However, for pre-quench ground states in the paramagnetic phase, the numerical results are partly inconclusive.
Fig. 3 shows a starting point at $\Delta_{i}=-0.5$ successively approaching the quadruple point. When the quench starts far from this multiphase point, the numerical results support the conjecture. However, as the quench starting point moves closer to the quadruple point, violations of the conjecture begin to appear close to the Ising phase transition. Fig. 5 explores whether these violations could be caused by finite-size effects, as a shifted position of the phase boundaries for smaller systems is a known phenomenon in the ANNNI model [17].
@@
[…]
Figure 3: Quenches within the paramagnetic phase of starting point successively approaching the quadruple point.
The concept of a “phase transition” is not well-defined in finite systems, particularly when the considered systems are very small. The term only applies to systems in the thermodynamic limit, where a phase transition is commonly characterised by a discontinuity in an order parameter. In contrast, at finite sizes, such discontinuities cannot be found, as it broadens to be a smooth extremum of the order parameter.
However, a quantity known as fidelity susceptibility has been shown to serve as an effective indicator of quantum phase transitions [22, 23, 24] as it quantifies the sensitivity of the ground state to perturbations. The fidelity susceptibility $\chi$ can be derived from Taylor expanding the fidelity $\mathcal{F}$ that is defined as [24]
$\mathcal{F}(a,\delta a)=|\left\langle\text{GS}(a)\mid\text{GS}(a+\delta a)\right\rangle|,$ (48)
where $a$ denotes a parameter of the Hamiltonian and $\delta a$ a small change in $a$. Its Taylor expansion reads [22]
$\mathcal{F}(\delta a)=1-\frac{1}{2}\chi(\delta a)^{2}+\mathcal{O}\left((\delta a)^{3}\right).$ (49)
The second order expansion coefficient sets the fidelity susceptibility $\chi$ to be
$\chi=\frac{2}{(\delta a)^{2}}(1-\mathcal{F}).$ (50)
@@
The location of the phase transition found via this method identified using this method generally approaches the infinite-size phase transition asymptotically. The technique is applied in two directions, given that the phase diagram is two-dimensional: the shift in $h$ direction was determined while keeping $\Delta$ constant, and vice versa.
The “remodelled” phase boundaries for a system of $L=16$ spins are depicted in Fig. 6. We are aware that the fidelity susceptibility is not a reliable measure for tracking phase transitions near the floating phase, as it typically exhibits peak revivals that depend on the system size (cf. [25]).
Some violations, such as those illustrated in Fig.s 3 and 4, remain inexplicable by the adjusted positions of the phase transitions. However, other instances, such
@@
as the quenches depicted in Fig.s 4, 5, and 5, can be accounted for by these revised phase transition locations.
VI Overlaps in Post-Quench Spectrum
For quenches within the same phase, the conjecture states that the largest overlap with the initial ground state should be that with the final ground state. Consequently, when plotting the overlaps against the post-quench Hamiltonian spectrum, the largest peak should align with the lowest energy state. In the following plots, we shifted the post-quench spectrum for this purpose such that the ground state energy is located at $0.0$ and all excitations energies are positive.
@@ VI Overlaps in Post-Quench Spectrum
Figure 6: Finite-size shifted phase boundaries of the ANNNI model when keeping the parameters $h$ (panel 6) and $\Delta$ (panel 6) fixed, respectively. The dashed lines mark the DMRG-calculated phase boundaries of the infinite-sized system [14], while the red lines show the shifted transition lines.
Fig. 7 shows quenches within the paramagnetic phase, carried out along the depicted trajectories. The direction labelled “forward” denotes a quench from $i$ to $f$, while “backward” indicates a quench from $f$ to $i$. Quenches within the floating phase, anti-phase, and ferromagnetic phase are presented below using the same labelling scheme.
The results in Fig.s 7 to 10 were obtained employing exact diagonalisation. They support the conjecture, as the largest peak consistently occurs at the ground state energy.
This rendering of the results allows for a re-investigation of the quench depicted in Fig. 4, which fails to support the conjecture as points near the Ising transition exhibit a larger overlap than the ground state overlap. A parameter combination for the quench starting at the same point as in Fig. 4 and ending in the darker region close to the phase boundary is used for the result illustrated in Fig. 11.
@@
Figure 7: Quenches within the paramagnetic phase. The phase diagram in panel 7 shows the starting and ending points of the quenches. The quench from $i$ to $f$ is referred to as “forward”. Adapted from Karrasch et al. [14].
@@
Figure 8: Quenches within the ferromagnetic phase. The phase diagram in panel 8 shows the starting and ending points of the quenches. The quench from $i$ to $f$ is referred to as “forward”. Adapted from Karrasch et al. [14].
@@
Figure 9: Quench within the anti-phase. The phase diagram in panel 9 shows the starting and ending points of the quench. The quench from $i$ to $f$ is referred to as “forward”. Adapted from Karrasch et al. [14].

### v2

@@ IV Axial Next Nearest Neighbor Ising Model
[…]
Figure 2: Quenches within the (a) ferromagnetic and (b) floating phases and (c) the antiphase.
@@
The competition between interactions gives rise to a more complex phase diagram, accommodating four different phases, as illustrated in Fig. 1. These phases are the paramagnetic (PM) phase, the ferromagnetic (FM) phase, a critical incommensurate floating phase (FP), and an antiphase (AP).
In contrast to the TFIM, the ANNNI model is not integrable [32], except along two lines in the phase diagram: the trivial line with $\Delta=0$ (which corresponds to the TFIM) and the so-called Peschel–Emery (PE) line [36]. The PE line is a disorder line along which the Hamiltonian factorizes into local Hamiltonians $H_{j}^{\operatorname{loc}}$, such that it takes the form [37]
$H_{\mathrm{ANNNI}}^{\mathrm{PE}}=\sum_{j}H_{j}^{\operatorname{loc}}.$ (37)
The ground state of this factorized Hamiltonian is exactly degenerate and can be expressed as a product state [38]. The model described by Eq. (37) is frustration-free, as the ground state of the full Hamiltonian $H_{\mathrm{ANNNI}}^{\mathrm{PE}}$ minimizes each local Hamiltonian $H_{j}^{\operatorname{loc}}$ independently [37]. In the ANNNI model, the PE line is described by [36]
@@
Figure 3: Quenches within the paramagnetic phase of the starting point successively approaching the quadruple point.
The Hamiltonian of the ANNNI model can be mapped onto an interacting quantum Ising model by dual mapping and a subsequent rotation [38]. After the further application of a Jordan–Wigner transformation, this Hamiltonian is equivalent to the Kitaev–Hubbard chain [38, 37, 39]:
$H_{\mathrm{ANNNI}}$ $=J\sum_{j=1}^{L}\big[h(c_{j}^{\dagger}-c_{j})(c_{j+1}^{\dagger}+c_{j+1})$ (39)
$+\Delta(1-2c_{j}^{\dagger}c_{j})(1-2c_{j+1}^{\dagger}c_{j+1})-(1-2c_{j}^{\dagger}c_{j})\big].$
Due to the quartic interaction arising from the next-nearest neighbor coupling in the spin-$\frac{1}{2}$ system, the ANNNI model does not have a simple analytical solution like the TFIM.
As the Hamiltonian factorizes, the ground state along the PE line is a direct product state [37]:
$|\text{GS}_{i}^{(\pm)}\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}\big(\left|\uparrow\right\rangle_{j}\pm\alpha\left|\downarrow\right\rangle_{j}\big),$ (40)
where $\alpha=\sqrt{\cot{\left(\frac{\theta^{*}}{2}\right)}}$} with ${\theta^{*}=\arctan{\left(h^{*}\right)}}=\arctan{\left(\Delta-\frac{1}{4\Delta}\right)}$ [37]. This ground state could be normalized by the factor $(1+\alpha^{2})^{-L/2}$, but we omit normalization because it is not required for the following discussion. We see that $\alpha\geq 1$ along the entire frustration-free line.
@@
[…]
@@
[…]
$\hskip-5.69046pt|\psi^{(\pm)}\rangle=\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{N}}\left(|\uparrow\rangle_{j}\pm\alpha|\downarrow\rangle_{j}\right)\mathbin{\mathop{\otimes}\displaylimits_{l=N+1}^{L}}\left(\pm\alpha|\uparrow\rangle_{l}-|\downarrow\rangle_{l}\right)$ (41)
for $N<L$. These states, like the ground state, are not normalized. Note, that these states are exact eigenstates only along the PE line.
The overlap between two ground states along the PE line with positions characterized by $\alpha_{i}$ and $\alpha_{f}$, respectively, then reads
$\big|\langle\text{GS}_{i}^{(\pm)}$ $|\text{GS}_{f}^{(\pm)}\rangle\big|^{2}$
$=$ $\ \bigg|\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}\left(\left\langle\uparrow\right|_{j}\pm\alpha_{i}\left\langle\downarrow\right|_{j}\right)\mathbin{\mathop{\otimes}\displaylimits_{l=1}^{L}}\left(\left|\uparrow\right\rangle_{l}\pm\alpha_{f}\left|\downarrow\right\rangle_{l}\right)\bigg|^{2}$
$=$ $\ \big|(1+\alpha_{i}\alpha_{f})^{L}\big|^{2}=(1+\alpha_{i}\alpha_{f})^{2L},$ (42)
and the overlap between a ground state and an excited state yields
$\big|\langle\text{GS}_{i}^{(\pm)}|\psi_{f}^{(\pm)}\rangle\big|^{2}=\bigg|$ $\mathbin{\mathop{\otimes}\displaylimits_{j=1}^{L}}\left(\left\langle\uparrow\right|_{j}\pm\alpha\left\langle\downarrow\right|_{j}\right)\mathbin{\mathop{\otimes}\displaylimits_{l=1}^{N}}\left(|\uparrow\rangle_{l}\pm\alpha|\downarrow\rangle_{l}\right)$
$\hskip-4.26773pt\times\mathbin{\mathop{\otimes}\displaylimits_{m=N+1}^{L}}\left(\pm\alpha|\uparrow\rangle_{m}-|\downarrow\rangle_{m}\right)\bigg|^{2}$ (43)
$=$ $\ (1+\alpha_{i}\alpha_{f})^{2N}(\pm\alpha_{f}\mp\alpha_{i})^{2(L-N)}.$
$\hskip-4.26773pt\big|\langle\text{GS}_{i}|\text{GS}_{f}\rangle\big|^{2}\overset{!}{>}\big|\langle\text{GS}_{i}|\psi_{f}\rangle\big|^{2}\Leftrightarrow\Bigg|\frac{\langle\text{GS}_{i}|\psi_{f}\rangle}{\langle\text{GS}_{i}|\text{GS}_{f}\rangle}\Bigg|^{2}\overset{!}{<}1.$ (44)
@@
Figure 6: Finite-size shifted phase boundaries of the ANNNI model when keeping the parameters (a) $h$ and (b) $\Delta$ fixed, respectively. The dashed lines mark the DMRG-calculated phase boundaries of the infinite-sized system [32], while the red lines show the shifted transition lines.
@@
[…]
Figure 7: Quenches within the paramagnetic phase. (a) Phase diagram indicating the starting $i$ and ending $f$ points of the quenches. (b), (c) Magnitude of the overlap between the prequench ground state at $i$ and each eigenstate of the finite system at $f$, shown for the vertical (b) and horizontal (c) quenches marked in (a). A quench from $i$ to $f$ is referred to as “forward.” Adapted from Karrasch and Schuricht [32].
Calculating this ratio gives
$\bigg|\frac{\langle\text{GS}_{i}|\psi_{f}\rangle}{\langle\text{GS}_{i}|\text{GS}_{f}\rangle}\bigg|^{2}$ $=\frac{(1\pm\alpha_{i}\alpha_{f})^{2N}(\pm\alpha_{f}\mp\alpha_{i})^{2(L-N)}}{(1+\alpha_{i}\alpha_{f})^{2L}}$
$=\bigg(\frac{\pm\alpha_{f}\mp\alpha_{i}}{1+\alpha_{i}\alpha_{f}}\bigg)^{2(L-N)}.$ (45)
To decide whether this expression is larger or smaller than 1, we recall that $\alpha\in[1,\infty)$. Thus, we can distinguish four cases.
  1. (1)
    $\alpha_{i}\rightarrow 1$, and $\alpha_{f}\rightarrow\infty$.
  2. (2)
    $\alpha_{i}\rightarrow\infty$, and $\alpha_{f}\rightarrow 1$.
  3. (3)
    $\alpha_{i}\rightarrow 1$, and $\alpha_{f}\rightarrow 1$.
  4. (4)
    $\alpha_{i}\rightarrow\infty$, and $\alpha_{f}\rightarrow\infty$.
The limit for all four cases tends to zero and is thus smaller than 1. Therefore, the conjecture that can be rephrased into condition (44) is fulfilled.
The two product states $|\text{GS}^{(+)}\rangle$ and $|\text{GS}^{(-)}\rangle$ are both exact ground states and are exactly degenerate. Further, they are related by the global $\mathbb{Z}_{2}$ spin-flip symmetry but they belong to different symmetry sectors. Thus, their respective excitations $|\psi^{(+)}\rangle$ and $|\psi^{(-)}\rangle$ also belong to different sectors. As a consequence, the physically relevant overlaps are those within the same sector (i.e., between states carrying the same superscript).
[…]
Finally, Eq. (45) confirms that condition (44) is satisfied along the entire frustration-free line. Thus, the conjecture can be verified analytically for the special case of a quench along this line.
@@
[…]
Figure 8: Quenches within the ferromagnetic phase. The phase diagram in (a) shows the starting and ending points of the quenches. The quench from $i$ to $f$ is referred to as “forward.” Adapted from Karrasch and Schuricht [32].
@@
Figure 9: Quench within the antiphase. The phase diagram in (a) shows the starting and ending points of the quench. The quench from $i$ to $f$ is referred to as “forward.” Adapted from Karrasch and Schuricht [32].
V Numerical Analysis of the ANNNI model
To investigate the conjecture’s validity beyond this special case, we conducted a numerical analysis of the ANNNI model. We present results obtained through exact diagonalisation (ED). Given the constraints of the available computational resources, the largest accessible system size for these computations is $L=16$ spins. The presentation of the results is structured according to the phase diagram shown in Fig. 1. Each point in the phase diagram corresponds to a parameter combination that uniquely defines the postquench Hamiltonian. For clarity and ease of interpretation, dashed lines are used to reconstruct the phase boundaries. The quench starting points are indicated above each plot and marked by a star in the phase diagram. The results are displayed using a color code based on the normalized deviation of the ground state overlap from the maximal overlap, defined as
$\text{deviation}=\frac{|\text{GS}\rangle\ \text{overlap}-\text{max overlap}}{\text{max. overlap}}.$ (46)
In this visualization, a very bright teal color indicates that the ground state overlap equals the maximal overlap, while darker colors signify a greater deviation between these two quantities.